OpenStudy (anonymous):

WILL GIVE MEDAL: solve tan(5pi/3). unit circle question, please help. the answer is - root 3, how do you get that?

OpenStudy (anonymous):

ok do you have your circle in front of you?

OpenStudy (anonymous):

i don't use a unit circle, i use the 2 triangles

OpenStudy (anonymous):

i somehow did it before like that?

OpenStudy (anonymous):

If you look at your circle 5\[ \frac{ 5\pi }{ 3} \] is equal to having the coordinates of \[\frac{ 1 }{ 2 }\] \[-\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

and you know that tan is b/a so than you get \[-\sqrt{3}\]

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

hmm ok thx

OpenStudy (triciaal):

this is a bit longer but it shows more an angle of 5pi/3 is the angle of 5pi/3*180/pi = 300 degrees for a circle same as -60 degrees tan 300 = -tan 60 = -1.732 for a rt triangle 30-60-90 the measure of the sides are in the ratio

OpenStudy (anonymous):

ohhh thanks that helped a lot better, :)